# PRISM-Consult: A Panel-of-Experts Architecture for Clinician-Aligned Diagnosis

### v1

@@ Abstract
We present PRISM-Consult, a clinician-aligned panel-of-experts architecture that extends the compact PRISM sequence model into a routed family of domain specialists. Episodes are tokenized as structured clinical events; a light-weight router reads the first few tokens and dispatches to specialist models (Cardiac–Vascular, Pulmonary, Gastro–Oesophageal, Musculoskeletal, Psychogenic). Each specialist inherits PRISM’s small transformer backbone and token template, enabling parameter efficiency and interpretability. On real-world Emergency Department cohorts, specialists exhibit smooth convergence with low development perplexities across domains, while the router achieves high routing quality and large compute savings versus consult-all under a safety-first policy. We detail the data methodology (initial vs. conclusive ICD-9 families), routing thresholds and calibration, and report per-domain results to avoid dominance by common events. The framework provides a practical path to safe, auditable, and low-latency consult at scale, and we outline validation steps—external/temporal replication, asymmetric life-threat thresholds, and multi-label arbitration—to meet prospective clinical deployment standards.
@@ II Introduction
II-A Why a Panel of Experts? Tokenization Trade-offs and Clinical Coverage
@@
This creates a fundamental tension in single-model designs: optimizing the token schema for one domain (e.g., cardiology) may degrade performance in another (e.g., pulmonary embolism vs. pneumothorax), where different signals and thresholds matter. PRISM-Consult resolves this by holding the token template constant—preserving interpretability and shared embeddings—while routing episodes to domain-specialized fine-tunings. Each specialist selectively expands or emphasizes a local subset of tokens relevant to its domain (e.g., pleural processes, gas exchange markers for pulmonary; ischemic markers and ECG patterns for cardiac), without forcing the global vocabulary to balloon. The router thus acts as a clinical analogue of referral: “right expert, right time,” mitigating the specificity–coverage trade-off inherent to a single global model.
@@ III Related Work
III-A Clinical foundation models and domain adaptation
@@
III-B Mixture-of-experts and routing among specialists
@@
III-C Positioning PRISM-Consult
@@ IV-B1 Data Source
Data for this study was extracted from data from the MIMIC-IV database, an extensive electronic health records repository collected from patients admitted to the Beth Israel Deaconess Medical Center (BIDMC) between 2008 and 2019. The database encompasses detailed clinical information, including demographics, vital signs, laboratory test results, diagnostic procedures, and discharge diagnoses captured during hospital stays [19].
IV-B2 Data structure and cohort window.
@@
IV-B3 Inclusion/exclusion Criteria.
@@ IV-B5 Final (Conclusive) ICD-9-CM Code Set — Specialist Gold Labels
00footnotetext: Per prior protocol, include locally used acute pericarditis/myocarditis codes in this family and harmonize under the same training label.
IV-B6 Sequence construction.
Once eleigible patients were determined, their medical histories were extracted and converted to token sequences using a fixed event-precedence policy, based foremost on chronological timing and then a predetermined hierarchy of token types: DIAG >> LAB >> ORDER, with explicit time-gap markers inserted between events when inter-event intervals exceed pre-specified thresholds (e.g., 1 hour, 6 hours) to preserve chronology. A final alphabetical ordering within each token-type is then done to ensure that batched sequences are ordered consistently across episodes (this ensures consistency in prediction accuracy given the singular ’next token’ generation schema to the PRISM model). Each sequence is truncated or windowed to a maximum length of 512 tokens for an entire patient chronology, potentially spanning multiple admission episodes. The final (gold-label) diagnosis token is not included on the input side for training tasks that predict it, preventing label leakage.
IV-B7 Tokenization template and vocabulary.
@@ IV-C2 Evaluation (PRISM baseline)
IV-D PRISM-Consult (panel of experts) Confguration and Training
@@
    A router trained on the earliest events of each episode to emit one or more domain flags (Cardiac–Vascular, Pulmonary, Gastro–Oesophageal, Musculoskeletal, Psychogenic). Inputs are the first 2–5 coded events represented as bag-of-concepts TF–IDF features projected to a 256-dim space, optionally concatenated with simple temporal features (e.g., time since triage). The router is a 2-layer transformer classifier with sigmoid outputs and focal loss to up-weight life-threatening domains.
@@ IV-D2 Calibration, safety, and auditability
IV-D3 Compute and reproducibility.
Experiments for the router were run on modern GPUs (e.g., A100-class) using deterministic seeds. Data preprocessing and training pipelines are version-controlled; model checkpoints, tokenizer vocabularies, and ICD-9 codebooks (both initial and final sets) are archived with SHA checksums.
@@ IV-D9 Router model and calibration
Compute proxy and latency.
@@ V-A2 Domain-specific outcomes (Specialist Model Set Performance)
Cardiac–Vascular (AMI/UA/pericardial/aortic/HCM).
@@
Pulmonary (PE/pneumothorax/pleura/pneumonia).
@@
Gastro–Oesophageal (GERD/spasm/Boerhaave/ulcer perforation).
@@
Musculoskeletal (costochondritis/rib injury/chest-wall contusion).
@@
Psychogenic (panic/anxiety/hyperventilation).
@@ VI-A Cross-disciplinary performance of PRISM
VI-B Next steps: validation and extension
@@
VI-C Limits
Three limitations temper interpretation. First, the present corpus is effectively single-label across directories, so the multi-label routing regime was not strongly exercised; future cohorts should include confirmed multi-diagnosis cases to probe arbitration. Second, class imbalance (e.g., the relatively small Musculoskeletal pool) likely contributes to higher perplexity and calibration variance for that head; targeted sampling, modest adapter rank increases, and data augmentation (token normalization for near-synonymous events) are straightforward mitigations. Third, learned absolute positional embeddings fix the maximum context length and may not extrapolate; if longer horizons are needed, rotary or ALiBi encodings can be substituted with minimal disruption to the rest of the stack. Finally, threshold selection in the current run favored a near-constraint point (life-threat recall <0.98<0.98 on test); future reports should hard-enforce prespecified safety constraints and present confidence intervals via bootstrap over episodes.

### v2

@@ Abstract
We present PRISM-Consult, a clinician-aligned panel-of-experts architecture that extends the compact PRISM sequence model into a routed family of domain specialists. Episodes are tokenized as structured clinical events; a light-weight router reads the first few tokens and dispatches to specialist models (Cardiac–Vascular, Pulmonary, Gastro–Oesophageal, Musculoskeletal, Psychogenic). Each specialist inherits PRISM’s small transformer backbone and token template, enabling parameter efficiency and interpretability. This initial study evaluates a scoped panel of five specialist families defined by high-impact ED diagnostic groups. On real-world Emergency Department cohorts, specialists exhibit smooth convergence with low development perplexities across domains, while the router achieves high routing quality and large compute savings versus consult-all under a safety-first policy. We detail the data methodology (initial vs. conclusive ICD-9 families), routing thresholds and calibration, and report per-domain results to avoid dominance by common events. The framework provides a practical path to safe, auditable, and low-latency consult at scale, and we outline validation steps—external/temporal replication, asymmetric life-threat thresholds, and multi-label arbitration—to meet prospective clinical deployment standards.
@@ II INTRODUCTION
II-A Clinical motivation and use case
In the Emergency Department (ED), clinicians must make time-critical decisions from sparse, noisy early signals: a chief complaint, a small number of initial vitals, and the first diagnostic actions (e.g., ECG, labs, imaging orders). At this stage, the task is rarely “predict the final diagnosis” in a single step; instead, clinicians form an evolving differential and decide which actions are most urgent and informative. This creates a practical need for early decision support that can suggest likely next steps and plausible differentials while remaining fast, auditable, and conservative under uncertainty.
II-B Why clinician-aligned routing
Real ED care is organized around consult services and organ-system expertise (e.g., cardiology, pulmonary/critical care, gastroenterology). We therefore frame early prediction as a routing problem: from the earliest events, decide which specialist perspectives are most appropriate to engage. This design supports safety-first behavior (e.g., consult multiple specialists when uncertain) and enables explicit control over compute/latency trade-offs while preserving clinical interpretability: a routed specialist corresponds to a clinically recognizable “who should weigh in?” decision.
II-C Why transformer specialists rather than a single monolith
A single sequence model can be effective, but ED trajectories often cross domains and contain heterogeneous patterns (e.g., dyspnea spanning pulmonary embolism, pneumonia, heart failure, anxiety). Specialization allows each model to learn domain-consistent thresholds, priors, and event semantics while retaining a shared tokenization template and backbone for parameter efficiency. The resulting panel improves coverage without requiring an ever-expanding single-model vocabulary, and it allows domain-specific calibration and evaluation aligned to clinical consequences.
@@ II-C Why transformer specialists rather than a single monolith
II-D Why a Panel of Experts? Tokenization Trade-offs and Clinical Coverage
@@
This creates a fundamental tension in single-model designs: optimizing the token schema for one domain (e.g., cardiology) may degrade performance in another (e.g., pulmonary embolism vs. pneumothorax), where different signals and thresholds matter.
II-E Heterogeneity, Comorbidity, and Cross-Specialty Trajectories in the ED
Further complicating diagnostic predictions is the fact that emergency presentations are intrinsically heterogeneous and frequently comorbid: a single complaint (e.g., chest pain, dyspnea, syncope) often admits plausible explanations across multiple organ systems, and early measurements (vitals, first labs, first orders) carry overlapping, sometimes contradictory signals. In MIMIC-IV ED cohorts, the same prefix of events can legitimately evolve toward cardiac (AMI, dissection), pulmonary (PE, pneumothorax, pneumonia), gastro–oesophageal (Boerhaave, spasm, reflux), musculoskeletal, or psychogenic end points, or some combination thereof, and patients routinely carry chronic conditions (CKD, COPD, diabetes, atrial fibrillation) that shift priors and confound intermediate tests. These realities create diagnostic pathways that traverse traditional service boundaries and unfold via cross-specialty handoffs.
This cross-disciplinary drift introduces several modeling hazards for single-domain specialists. First, boundary undercoverage: a specialist tuned to one domain’s vocabulary and thresholds may systematically down-weight rare but life-threatening out-of-domain hypotheses that share early features (e.g., PE vs. AMI in troponin-positive dyspnea). Second, calibration shift: domain-specific likelihoods calibrated in-domain can become miscalibrated when the latent etiology changes mid-episode, especially under sparse, prefix-only supervision. Third, premature closure: strong local patterns (e.g., ECG orders $\rightarrow$ ischemia) can anchor the trajectory before counter-signals appear (e.g., pleuritic descriptors, D-dimer, CTA), yielding myopic next-event proposals.
Sequential structure further compounds these risks. Clinical events arrive as partially ordered sets (co-occurring orders, batched lab panels) with irregular time gaps, and many ED timelines are better viewed as sequences of sets than pure token strings. Methods that enforce set-level invariance or augment code streams with ontology signal (e.g., CCS/knowledge-graph embeddings) help preserve meaning across co-occurrences and synonyms, but they do not by themselves resolve cross-service drift when the generator of observations changes during workup[32] [33]. Likewise, zero-shot generative rollouts over patient-health timelines demonstrate breadth but still reflect the chosen pretraining mix and may blur service-specific decision thresholds [29].
From a probabilistic standpoint, the ED setting resembles large-context inference over many interacting latent causes. Classic graphical approaches (e.g., QMR-DT) formalize comorbidity and conditional dependence but become computationally brittle at realistic scales [34]. Autoregressive transformers approximate these high-order conditionals efficiently, yet as already noted above, a single backbone must compromise between compact vocabularies (speed, stability) and domain-specific expressivity (rare tokens, narrow thresholds). In practice, optimizing tokenization and priors for one service can degrade discrimination in another, particularly near safety-critical boundaries.
These observations motivate a routed, multi-disciplinary engagement approach. Concretely it: (i) reads the earliest prefix and estimate calibrated probabilities over clinical domains; (ii) consults one or more domain specialists when warranted (with asymmetric, safety-first thresholds for life-threatening etiologies); (iii) arbitrates multi-label suggestions with deterministic, auditable policies; and (iv) fails open when confidence is low or vitals breach guardrails. Such conditional computation aligns with mixture-of-experts and expert-routing advances—activating only the parameters relevant to the emerging hypothesis space while preserving latency and traceability.
PRISM-Consult is designed with this approach in mind. By holding the token template constant—preserving interpretability and shared embeddings—while routing episodes to domain-specialized fine-tunings. Each specialist selectively expands or emphasizes a local subset of tokens relevant to its domain (e.g., pleural processes, gas exchange markers for pulmonary; ischemic markers and ECG patterns for cardiac), without forcing the global vocabulary to balloon. The router thus acts as a clinical analogue of referral: “right expert, right time,” mitigating the specificity–coverage trade-off inherent to a single global model.
@@ III Related Work
III-A Diagnostic Progressions and LLM Token Prediction
In large language models (LLMs), each token—representing a word or sub-word unit—has clear pairwise probabilities relative to other tokens. For example, given a token such as ”New,” the subsequent token might have high probability for ”York,” ”Zealand,” or ”Jersey.” Individually, these pairwise probabilities are straightforward. However, as the sequence of tokens expands—considering an entire preceding sentence or paragraph—the combinatorial complexity and contextual dependencies make it impossible to precisely calculate probabilities using first order correlations alone. Hence, deep learning and attention mechanisms in transformer architectures (like GPT) approximate these complex conditional dependencies, capturing long-range contexts and intricate relationships between tokens.
By analogy, diagnostic medicine faces a similar complexity. Individual diagnostic tests may exhibit well-defined pairwise probabilities for specific diseases—just as pairwise token probabilities can be well-established. However, when attempting to predict the next best test or disease diagnosis given the entire preceding sequence of events (the patient’s complete history, including medical tests, symptoms, prior diagnoses, lifestyle factors, and comorbidities), the scenario becomes profoundly more complex. The interactions between these elements are intricate, context-dependent, and frequently nonlinear, making exact probabilistic inference practically infeasible.
Thus, Bayesian statistical approaches quickly reach limitations. Transformer-based diagnostic models, analogous to language models, can approximate the complex contextual dependencies inherent in patient histories, making them potentially more capable of accurately predicting subsequent medical events (tests, results, or diagnoses). These models implicitly learn from data to reflect patient-level complexity more naturally, offering both predictive capability and interpretable attention scores to highlight influential factors in diagnostic reasoning.
III-B Prior Work in Transformer and Deep Learning Models for Sequential Clinical Diagnosis
In contrast to high-level summations of pre-existing efforts, which are considerable and highly varied in nature, we opt to focus in on prior work of immediate relevance to our proposed approach, by zeroing in on a methodology-specific approaches that have informed our own experimental design, and serve as a basis for which our work attempts to build on.
III-B1 Tokenization and Representation of Medical Data
The first step in adapting transformer models to diagnostic-sequence prediction is to convert a patient’s richly structured history into a sequence of discrete tokens. Early work relied on standardised vocabularies such as ICD-9/ICD-10, CPT and Read codes, treating each code as an individual token [30, 31]. A coarser alternative aggregates thousands of raw codes into a few hundred clinically coherent groups, reducing sequence length while retaining clinical meaning [30].
More granular pipelines tokenise free-text concepts, symptoms and history fragments extracted from clinical narratives [28]. The ETHOS framework extend this idea further by encompassing: admissions, diagnoses (ICD-10-CM), procedures (ICD-10-PCS), medications (ATC), laboratory results (LOINC) and even inter-event time gaps. Under the ETHOS framework, these events are mapped to 1–7 tokens each, then ordered chronologically into a patient-health timeline (PHT) [29].
Because multiple events can occur simultaneously during a visit, several authors represent a patient record as a sequence of sets, forcing the model to be order-invariant within each set. DPSS is the canonical example of this strategy [32]. Finally, semantic enrichment with external knowledge graphs (or hierarchical ontologies such as CCS) can be injected by concatenating concept embeddings with ontology embeddings, as done in the SETOR framework [33].
III-B2 Transformer Model Architectures in Diagnostic Prediction
Traditionally model choice varied depending on the prediction task. Encoder-only families (BERT, ALBERT, MedAlbert) excel at classification problems such as early disease detection from longitudinal code streams; MedAlbert, for instance, uses a 6-layer ALBERT encoder to flag incipient lung cancer three years in advance [30]. Decoder-only designs (GPT-like) suit generative forecasting: ETHOS, for instance, employs a causal decoder to roll out future PHTs in a zero-shot setting [29]. Hybrid encoder–decoder stacks (e.g., TransformEHR) translate past encounters into predicted future ICD sequences [31].
Domain-aware modifications are common. SETOR feeds visit-level embeddings— augmented with CCS-derived ontology vectors and continuous-time positional encodings—into a multi-layer transformer encoder dubbed the Patient Journey Transformer [33]. Such customisations can improve both performance and interpretability without abandoning the core attention machinery of the original architecture.
III-C Additional Clinical foundation models and domain adaptation
@@
III-D Mixture-of-experts and routing among specialists
@@
III-E Clinical sequence similarity and pathway clustering
A long line of work in clinical sequential pattern mining and pathway analysis focuses on defining similarity between patient trajectories and clustering them into clinically meaningful pathways. Temporal similarity measures are central in this setting because clinical workflows depend not only on event identity but also on ordering, repetition, and time gaps [36]. Pathway clustering methods adapt classical alignment/edit-distance algorithms (e.g., Needleman–Wunsch variants) to group patient journeys into interpretable clusters [37]. Related studies also examine clustering diagnosis-code sequences jointly with demographics to discover stratified pathway structure [39]. These approaches primarily target cohort discovery, summarization, and retrospective pathway characterization; by contrast, PRISM-Consult targets prospective early-prefix prediction and clinician-aligned dispatch to specialist models. Nonetheless, similarity-based pathway work informs what constitutes a “near miss” versus a “severe miss,” motivating our cost-sensitive evaluation.
III-F Probabilistic sequence models and stochastic automata
Classical probabilistic approaches such as Markov models, hidden Markov models, and stochastic automata have been applied to encoded clinical sequences due to their interpretability and tractability [41]. These methods can perform well when the state space is limited and the temporal process is relatively stationary. However, ED event streams are heterogeneous and high-dimensional, with long-range dependencies and sparse early prefixes that are difficult to capture with low-order Markov assumptions without aggressive state aggregation. We therefore focus on compact transformer specialists with shared tokenization, while explicitly comparing against and discussing probabilistic alternatives where appropriate.
III-G Positioning PRISM-Consult
@@ IV-B1 Data Source
Data for this study was extracted from the MIMIC-IV database, an extensive electronic health records repository collected from patients admitted to the Beth Israel Deaconess Medical Center (BIDMC) between 2008 and 2019. The database encompasses detailed clinical information, including demographics, vital signs, laboratory test results, diagnostic procedures, and discharge diagnoses captured during hospital stays [19].
IV-B2 Data structure and cohort window
@@
IV-B3 Inclusion/exclusion Criteria
@@ IV-B5 Final (Conclusive) ICD-9-CM Code Set — Specialist Gold Labels
Cohort composition note: The cohort is defined by the study’s previously delineated gold-label diagnostic families rather than the unconditional ED case mix; as a result, domain prevalence reflects selection criteria and may over-represent certain high-acuity presentations (e.g., cardiopulmonary). To mitigate dominance by any single domain, we report per-domain results in addition to macro-averaged metrics, and we treat evaluation on broader, less-selected cohorts as an external validation objective.
IV-B6 Sequence construction
Once eligible patients were determined, their medical histories were extracted and converted to token sequences using a fixed event-precedence policy, based foremost on chronological timing and then a predetermined hierarchy of token types: DIAG >> LAB >> ORDER, with explicit time-gap markers inserted between events when inter-event intervals exceed pre-specified thresholds (e.g., 1 hour, 6 hours) to preserve chronology. A final alphabetical ordering within each token-type is then done to ensure that batched sequences are ordered consistently across episodes (this ensures consistency in prediction accuracy given the singular ’next token’ generation schema to the PRISM model). Each sequence is truncated or windowed to a maximum length of 512 tokens for an entire patient chronology, potentially spanning multiple admission episodes. The final (gold-label) diagnosis token is not included on the input side for training tasks that predict it, preventing label leakage.
IV-B7 Tokenization template and vocabulary
@@ IV-C2 Evaluation (PRISM baseline)
IV-D PRISM-Consult (panel of experts) Configuration and Training
Early ED prefix (first $K$ events) Tokenization + optional time features Lightweight calibrated router $p(d\mid\text{prefix})$ over domains Dispatch policy Top-1 / Top-2 specialists Fail-open under low confidence Safety guardrails: low confidence or danger thresholds $\Rightarrow$ consult-all Domain specialists PRISM backbone + LoRA adapters Domain-aligned next-event / diagnosis-family scores Deterministic arbitration / merge (e.g., Cardiac >> Pulmonary >> …) Outputs Predictions + uncertainty Audit log: routed domains, scores, thresholds
Figure 3: PRISM-Consult end-to-end routing pipeline.
@@ IV-D PRISM-Consult (panel of experts) Configuration and Training
    A router trained on the earliest events of each episode to emit one or more domain flags (Cardiac–Vascular, Pulmonary, Gastro–Oesophageal, Musculoskeletal, Psychogenic). Inputs are the first 2–5 coded events represented as bag-of-concepts TF–IDF features projected to a 256-dim space, optionally concatenated with simple temporal features (e.g., time since triage). The router is a logistic-regression classifier and a Platt scaling router with sigmoid outputs and class weighting to up-weight life-threatening domains.
@@ IV-D2 Calibration, safety, and auditability
IV-D3 Compute and reproducibility
Experiments for the router were run on modern GPUs using deterministic seeds. Data preprocessing and training pipelines were version-controlled; model checkpoints, tokenizer vocabularies, and ICD-9 codebooks (both initial and final sets) are archived with SHA checksums.
Reproducibility and artifacts. All experiments were run on NVIDIA A100 (Ampere) GPU with 40GB VRAM, with with an 8-core CPU and 83.5 GB of system RAM) and software stack Linux Ubuntu, CUDA 12.x, NVIDIA proprietary driver.
We provide the full training and evaluation pipeline, including preprocessing scripts, tokenizer vocabulary construction, fixed random seeds, and configuration files. An accompanying public repository includes a minimal runnable notebook that reproduces key preprocessing and metric computations on a small illustrative example, enabling validation of the reported methodology.
Code availability. Code for PRISM-Consult, including preprocessing, training, and evaluation scripts, along with a minimal runnable example notebook, is available in the PRISM repository. [^3] https://github.com/lmlevine/prism
@@ IV-D9 Router model and calibration
Compute proxy and latency
@@ V-A2 Domain-specific outcomes (Specialist Model Set Performance)
Cardiac–Vascular (AMI/UA/pericardial/aortic/HCM)
@@
Pulmonary (PE/pneumothorax/pleura/pneumonia)
@@
Gastro–Oesophageal (GERD/spasm/Boerhaave/ulcer perforation)
@@
Musculoskeletal (costochondritis/rib injury/chest-wall contusion)
@@
Psychogenic (panic/anxiety/hyperventilation)
@@ V-D1 Safety and routing quality
Figure 8: End-to-end PRISM-Consult system summary (routing safety and efficiency)
@@ VI-A Cross-disciplinary performance of PRISM
VI-B Cost-sensitive clinical utility (future work)
While we report standard discrimination and ranking metrics, real clinical deployment is cost-sensitive: the consequence of a miss depends on severity (e.g., failing to identify a time-critical cardiopulmonary presentation) and the consequence of “over-consulting” is often primarily resource or latency cost. A principled extension is to report an expected-utility (or expected-cost) metric defined over the routed specialist set. For episode $i$ with true domain set $Y_{i}$ and routed set $R_{i}$, one can define:
$C_{i}=\sum_{d\in Y_{i}\setminus R_{i}}\lambda_{d}\;+\;\gamma\sum_{d\in R_{i}\setminus Y_{i}}1,$ (1)
where $\lambda_{d}$ is higher for life-threatening domains and $\gamma$ captures the comparatively lower cost of over-routing. This yields an explicit compute–safety Pareto curve as router thresholds vary. In the current study, we did not retain per-episode routed-set outputs for all runs, which prevents retrospective computation of such utility-based metrics from aggregate results alone. We therefore include this evaluation as a priority direction for future work and as a recommended standard for high-stakes clinical sequence prediction.
VI-C Next steps: validation and extension
@@
VI-D Limits
Three significant limitations temper interpretation:
Multi-label handling (limitation and evaluation plan)
First, the present corpus is effectively single-label across directories, so the multi-label routing regime was not strongly exercised. We therefore treat multi-label arbitration as a priority validation objective. In future work we will evaluate confirmed multi-diagnosis cases (e.g., AMI+PE) and report multi-label routing quality (e.g., Recallall@k / top-$k$ coverage, calibration by prefix length), along with structured audits of arbitration outcomes and fail-open behavior in safety-critical presentations.
Limited Clinical Domains and Skewed Patient-Mix
Second, this study evaluates only five specialist families selected to support clear gold-labeling and clinically meaningful ED coverage; this does not represent the full breadth of ED consult services. In addition, cohort composition reflects the selected diagnostic families and may be skewed toward higher-acuity presentations resulting in class imbalance (e.g., the relatively small Musculoskeletal pool) likely contributes to higher perplexity and calibration variance for that head; targeted sampling, modest adapter rank increases, and data augmentation (token normalization for near-synonymous events) are straightforward mitigations for class imbalance, and future work should also aim to expand the specialist panel (e.g., neurologic, infectious/sepsis, renal/metabolic, toxicologic/trauma) and evaluate scalability of routing and arbitration as the panel grows, and (ii) validate generalization on broader, less-selected cohorts and across sites/time periods to better reflect real-world ED case mix.
Fixed Positional Embeddings may Prove too Rigid for Model Adaptability
Third, learned absolute positional embeddings fix the maximum context length and may not extrapolate; if longer horizons are needed, rotary or ALiBi encodings can be substituted with minimal disruption to the rest of the stack.
Additionally, we note that threshold selection in the current run favored a near-constraint point (life-threat recall <0.98<0.98 on test); future reports should hard-enforce prespecified safety constraints and present confidence intervals via bootstrap over episodes.
Finally, to better contextualize model accuracy and efficiency, PRISM-Consult should be benchmarked against (i) classical classifiers (e.g., XGBoost) trained on the same early-prefix feature representation, and (ii) compact encoder models (e.g., ClinicalBERT/BioBERT-style) trained to predict domain/diagnosis from prefix tokens.